# Entanglement entropy bounds for pure states of rapid decorrelation

### v1

@@ 1 Introduction
A striking feature of the quantum entropy, in which it differs from its classical analog, is that in multicomponent systems the entropy of a state can be zero even when that of it’s restriction to a subsystem is strictly positive. In such cases the entropy of the reduced state is a measure of the quantum entanglement between the subsystem and the rest [44, 54]. One of our goals here is to present conditional decoupling criteria for an area-law type bound on the entanglement entropy of quantum states in general space dimensions.
@@ 1.1 Entanglement entropy
Our discussion is set in the context of quantum lattice systems whose single site components are qudits, whose state space is associated with the complex Hilbert space $\mathbb{C}^{\nu}$. The extended system’s Hilbert space is the tensor-product $\mathcal{H}_{W}:=\otimes_{u\in W}\mathbb{C}^{\nu}$, indexed by sites of a finite subset $W\subset\mathbb{Z}^{d}$. For each choice of a basis of $\mathbb{C}^{\nu}$, of $\nu$-elements whose collection we denote $\Omega_{0}$, there is a corresponding basis of $\mathcal{H}_{W}$ whose elements $|\boldsymbol{\sigma}\rangle$ are in $1$-$1$ correspondence with the collection of configurations, $\boldsymbol{\sigma}=(\sigma_{u})_{u\in W}\in\Omega_{0}^{W}=:\Omega_{W}$, of single-site variables with values in $\Omega_{0}$.
@@
at some $C<\infty$, which is independent of $A$, and a decoupling distance $l_{0}(A)$, whose dependence on $A$ relies on model-specific input.
@@ 1.2 Mutual information
where $S(A)$ is an abbreviation for the entropy of the reduced state $\varrho_{A}(\psi)$, and the last term is the conditional entropy $S(A_{1}\,|\,A_{2}):=S(A_{1}\cup A_{2})-S(A_{2})$. Viewed yet differently, the mutual information encodes the relative entropy of $\varrho_{A}(\psi)$ relative to the product state $\varrho_{A_{1}}(\psi)\otimes\varrho_{A_{2}}(\psi)$.
@@ 1.3 States of quasi-local structure in a computational basis
Our enabling criteria for entropy reduction refer to the structure of a state $|\psi\rangle\in\mathcal{H}_{W}$ as it appears in a convenient computational basis. Normalized pure states $|\psi\rangle$ are presentable in such a basis as
@@
in terms of a phase functions $\theta:\Omega_{W}\to(-\pi,\pi]$, $\theta(\boldsymbol{\sigma}):=\operatorname{Arg}\psi(\boldsymbol{\sigma})$ and a probability measure of weights $p(\boldsymbol{\sigma}):=|\psi(\boldsymbol{\sigma})|^{2}$ on $\Omega_{W}$.
Our discussion is not limited to but it simplifies in the case of states for which $\theta(\boldsymbol{\sigma})\equiv 0$. Such states are referred to as “sign-problem free” or stoquastic - a term invoking the relevance of a probabilistic perspective. Among the examples of stoquasticity are the ground states of positivity preserving Hamiltonians $H$. Included in that class are operators for which $(-H)$ in the given computational basis have only non-negative off-diagonal terms,
@@ 1.4 Example: the quantum Ising model
Among the states to which our general discussion applies are the ground states of the quantum Ising model in transversal field. The system’s Hilbert space is $\mathcal{H}_{W}=\otimes_{u\in W}\mathbb{C}^{2}$, and the Hamiltonian is
$H=-\sum_{u,v\in W}J_{u-v}\ S^{z}_{u}S^{z}_{v}-b\sum_{u\in W}S^{x}_{u}-h\sum_{u\in W}S^{z}_{u}\,,$ (1.10)
written in terms of elements of the Pauli triples of spin operators $\mathbf{S}_{u}=(S_{u}^{x},S_{u}^{y},S_{u}^{z})$ which act on the $u$-component of the tensor product. The coupling $J_{u-v}\geq 0$ is assumed to be finite range and ferromagnetic. A convenient computational basis which reveals the stoquasticity of $H$ is the eigenbasis of the local Pauli $S^{z}$-operators.
Restricted to $h=0$, the ground state undergoes a quantum phase transition at a threshold value $b_{c}$, which depends on $d$ and the couplings $J$. The transition is manifested in the structure of the infinite-volume limit of the ground state expectation value functional, of which a bellwether is the two point correlation function
$\langle S^{z}_{u}S^{z}_{v}\rangle^{(\mathbb{Z}^{d})}:=\lim_{L\to\infty}\langle S^{z}_{u}S^{z}_{v}\rangle^{(W_{L})}\,,$ (1.11)
where the superscript indicates the domain over with the model is formulated and $W_{L}=[-L,L]^{d}$. In these terms the model undergoes the following transition(cf. [48, 14]):
$\bullet$ For $b>b_{c}$, the ground state exhibits exponential decay of correlations:
@@
$\bullet$ For $b\in[0,b_{c})$ the ground state manifests long range order:
@@
$\bullet$ At $b=b_{c}$: $M(b)=0$ [13] but, unlike at $b>b_{c}$, the above spin-spin correlation decays by a dimension-dependent power law.
We shall not delve here into the direct formulation of an infinite-volume version of the model (cf. [17]), in which the (±)plus-or-minus($\pm$)( ± ) states are presentable as ground states of its infinite volume version. Instead, we shall focus on bounds for finite volumes $W_{L}$ which hold uniformly in $L$. Thus, unless specified explicitly the states refer to the model formulated in a finite subset $W\subset\mathbb{Z}^{D}$, which will be omitted from the notation. We also refrain here from the discussion of the effect of boundary conditions, which become of relevance for $b<b_{c}$.
The general results developed below, combined with known decay properties of the spin-spin correlations in this model [14, 23] enable us to prove the following.
@@
For the quantum Ising model on $W\subset\mathbb{Z}^{d}$, with the Hamiltonian (1.10) for any $b>b_{c}$:
@@
    The ground state is exponential clustering. In particular its truncated $x$-correlations decay exponentially: there is $C,\xi(b)\in(0,\infty)$ such that for any $W$, and $u,v\in W$:
@@
The bound (1.15) extends previous rigorous results on the entanglement in this model, which were restricted to one dimension and there to significantly sub-critical states  [33, 34]. For $d=1$ the logarithmic term in (1.15) is constant, yielding a strict area law. The bound on the mutual information and exponential clustering is a byproduct of our analysis. The bound (1.16) extends a result in [22] to the case $h=0$. It could potentially also be obtained using the probabilistic representation available for this model.
@@
2 State approximations of reduced complexity
A known approach towards bounds on the entanglement entropy $S(\varrho_{A}(\psi))$ in a specified pure state $|\psi\rangle$ is to seek high-fidelity and low-complexity approximations of its restriction to $A$ [10, 2, 8]. Seeking such approximations, we develop here a quantum version of conditioning on the state of a buffer $B$, of adjustable width $l$, enveloping the set $A$ and separating it from $C=W\setminus(A\sqcup B)$, cf. Figure 1.
@@ 2 State approximations of reduced complexity
Corresponding to the tripartition of $W$ we present the spin configurations as $\boldsymbol{\sigma}=(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\in\Omega_{A}\times\Omega_{B}\times\Omega_{C}$, and where convenient write the probability associated with $|\psi\rangle$ as
$|\psi(\boldsymbol{\sigma})|^{2}=p(\boldsymbol{\sigma})=p_{BC}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\ p_{A}(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\,,$ (2.1)
in terms of the probability of the variables in $B\sqcup C$, i.e. $p_{BC}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C}):=\sum_{\boldsymbol{\sigma}_{A}}p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})$, and the conditional probability $p_{A}(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})$. The subscripts on such probability functions will be omitted when deemed clear from the context, e.g. $p(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\equiv p_{BC}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})$, etc.
For a pure state $|\psi\rangle$ the reduced state in $A$ is the density matrix
@@
$\varrho_{A}(\psi)=$ $\sum_{\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C}}p(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\ |\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\rangle\langle\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})|$ (2.2)
with the normalized vector $|\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\rangle\in\mathcal{H}_{A}$, which is indexed by $(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\in\Omega_{B}\times\Omega_{C}$:
$|\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\rangle:=\sum_{\boldsymbol{\sigma}_{A}}e^{i\theta(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})}\sqrt{p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})}\ |\boldsymbol{\sigma}_{A}\rangle\,.$ (2.3)
Generically, the rank of the operator in (2.2) reaches $\min\left\{\dim\mathcal{H}_{A},\dim\mathcal{H}_{B\sqcup C}\right\}$. However, for states $|\psi\rangle$ for which the functions defining $|\psi_{A}(\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\rangle$ do not vary with $\boldsymbol{\sigma}_{C}$, the rank of $\varrho_{A}(\psi)$ drops to $\dim\mathcal{H}_{B}=\nu^{|B|}$.
Slightly more generally, the rank reduction holds for any state $|\psi\rangle$ for which the corresponding probability and phase functions satisfy the following pair of conditions:
$\mathrm{(I)}\qquad$ $p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})=p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B})\,,$

### v2

@@ 1 Introduction
A striking feature of the quantum entropy, in which it differs from its classical analog, is that in multicomponent systems, the entropy of a state can be zero even when that of its restriction to a subsystem is strictly positive. In such cases the entropy of the reduced state is a measure of the quantum entanglement between the subsystem and the rest [44, 54]. One of our goals here is to present conditional decoupling criteria for an area-law type bound on the entanglement entropy of quantum states in general space dimensions.
@@ 1.1 Entanglement entropy
Our discussion is set in the context of quantum lattice systems whose single site components are qudits, of state space associated with the complex Hilbert space $\mathbb{C}^{\nu}$. The extended system’s Hilbert space is the tensor-product $\mathcal{H}_{W}:=\otimes_{u\in W}\mathbb{C}^{\nu}$, indexed by sites of a finite subset $W\subset\mathbb{Z}^{d}$. For each choice of a basis of $\mathbb{C}^{\nu}$, of $\nu$-elements whose collection we denote $\Omega_{0}$, there is a corresponding orthonormal basis of $\mathcal{H}_{W}$. Its elements $|\boldsymbol{\sigma}\rangle$ are labeled by $\boldsymbol{\sigma}=(\sigma_{u})_{u\in W}\in\Omega_{0}^{W}=:\Omega_{W}$, configurations of single-site variables with values in $\Omega_{0}$
@@
at some $C<\infty$, which is independent of $A$, and a decoupling distance $l_{0}(A)$, whose dependence on $A$ relies on model-specific input. Section 6 elaborates on general conditions which ensure that the decoupling distance is bounded by a logarithm of the area, $l_{0}(A)\leq C\ln|\partial A|$.
@@ 1.2 Mutual information
where $S(A)\equiv S(\varrho_{A}(\psi))$ is an abbreviation for the entropy of the reduced state $\varrho_{A}(\psi)$, and the last term is the conditional entropy $S(A_{1}\,|\,A_{2}):=S(A_{1}\cup A_{2})-S(A_{2})$. Viewed yet differently, the mutual information encodes the relative entropy of $\varrho_{A}(\psi)$ relative to the product state $\varrho_{A_{1}}(\psi)\otimes\varrho_{A_{2}}(\psi)$.
@@ 1.3 States of quasi-local structure in a computational basis
The enabling criteria presented below for entropy reduction refer to the structure of a state $|\psi\rangle\in\mathcal{H}_{W}$ as it appears in a convenient computational basis in which it is presentable as
@@
in terms of a probability measure on $\Omega_{W}$ of weights $p(\boldsymbol{\sigma}):=|\psi(\boldsymbol{\sigma})|^{2}$, with rapid decay of correlations, and a phase function $\theta:\Omega_{W}\to(-\pi,\pi]$, $\theta(\boldsymbol{\sigma}):=\operatorname{Arg}\psi(\boldsymbol{\sigma})$ of similar properties.
The discussion is not limited to but it simplifies in the case of states for which $\theta(\boldsymbol{\sigma})\equiv 0$. Such states are referred to as “sign-problem free” or stoquastic - a term invoking the relevance of a probabilistic perspective. Among the examples of stoquasticity are the ground states of Hamiltonians $H$ for which $e^{-\beta H}$ is positivity-preserving in the specified computational basis. A sufficient condition for that is that the off-diagonal terms of $(-H)$ are real and non-negative.
@@ 1.4 Example: the quantum Ising model
Among the states to which our general discussion applies are the ground states of the quantum (or transversal field) Ising model (QIM). The system’s Hilbert space is $\mathcal{H}_{W}=\otimes_{u\in W}\mathbb{C}^{2}$, and the Hamiltonian (at $h^{z}\equiv 0$) is
$H=-\sum_{u,v\in W}J_{u-v}\ S^{z}_{u}S^{z}_{v}-b\sum_{u\in W}S^{x}_{u}\,,$ (1.10)
written in terms of elements of the Pauli triples of spin operators $(S_{u}^{x},S_{u}^{y},S_{u}^{z})$ which act on the $u$-component of the tensor product. The coupling $J_{u-v}\geq 0$ is assumed to be finite range and ferromagnetic. A convenient computational basis in which the model’s thermal and ground state are stoquastic is the eigenbasis of the local Pauli $S^{z}$-operators. Another such basis is the one in which the local $S^{x}$ are diagonal.
The system’s ground state undergoes a quantum phase transition at a threshold value $b_{c}>0$, which depends on $d$ and the couplings $J$. The transition is manifested in the behavior of the model’s two point correlation function, and most clearly in its infinite volume limit
$\langle S^{z}_{u}S^{z}_{v}\rangle^{(\mathbb{Z}^{d})}:=\lim_{L\to\infty}\langle S^{z}_{u}S^{z}_{v}\rangle^{(W_{L})}\,.$ (1.11)
The superscript on the correlation indicates the domain over which the model is formulated, with $W_{L}=[-L,L]^{d}\cap\mathbb{Z}^{d}$. In these terms, the model’s ground state undergoes the following transition (cf. [48, 14]):
$\bullet$ For $b>b_{c}$, its ground state correlation function decay exponentially:
@@
$\bullet$ For $b\in[0,b_{c})$ the correlation function exhibits long range order:
@@
$\bullet$ At $b=b_{c}$: $M(b)=0$ [13], but unlike at $b>b_{c}$ the spin-spin correlation decays by a dimension-dependent power law.
We shall not delve here into the direct formulation of the model’s infinite-volume version (cf. [17]), in which the (±)plus-or-minus($\pm$)( ± ) states are presentable as ground states of an infinite spin array. Instead, we shall focus on bounds for finite volumes $W_{L}$ which hold uniformly in $L$. Thus, unless specified explicitly, the states refer to the model formulated in a finite subset $W\subset\mathbb{Z}^{d}$, which will often be omitted in the notation.
We also refrain from the more complete discussion of boundary conditions. Unless specified otherwise there are to be assumed here as either free or constant (in terms of $\sigma^{z}$). For each of these, the Perron-Frobenius theorem applied to $\exp\left(-\beta H\right)$ in the $\sigma^{z}$-computational basis allows to conclude that the model’s finite volume ground states are unique at any $b\neq 0$.
The general results developed below, combined with known decay properties of the $S^{z}$-$S^{z}$-correlations in this model [14, 24] enable us to prove the following.
@@
For the quantum Ising model on $W\subset\mathbb{Z}^{d}$, $d\geq 1$, with the Hamiltonian (1.10) at $b>b_{c}(d)$ (i.e. throughout the ground states’ subcritical regime):
@@
    The ground state is exponentially clustering. In particular its truncated $x$-correlations decay exponentially: there is $C,\xi(b)\in(0,\infty)$ such that for any $W$, and $u,v\in W$:
@@
It should be noted that for $d=1$ where $\ln|\partial A|$ is constant, (1.15) yields a simple area law.
@@
Relations to previous results specific to QIM: The previous results on the entanglement specific to this model were restricted to one dimension and there to significantly sub-critical states [33, 34, 21]. The broader Theorem 1.1 addresses the entire subcritical regime, in all dimensions. The difference is traced below to a key handicap in the approach previously taken.
The bound on the mutual information and exponential clustering (1.16) extends a result of Crawford and Ioffe [23] that was derived using the model’s random current representation. It could potentially also be obtained using such methods.
2 State approximations of decreased complexity
A known approach towards bounds on the entanglement entropy $S(\varrho_{A}(\psi))$ of a specified pure state $|\psi\rangle$ is to seek high-fidelity approximations of the state $\varrho_{A}(\psi)$ in $A$ in terms of states of low-complexity, e.g. states of low rank – comparable with the exponential of the reduced state’s expected entropy [10, 2, 8]. Seeking such approximations, we present here two versions of approximations of $\varrho_{A}(\psi)$ invoking quantum versions of conditioning on the state in a buffer $B$ of adjustable width, enveloping the set $A$ and separating it from the much bigger $C=W\setminus(A\sqcup B)$, cf. Figure 1. The first is expressed directly in terms of the reduced state $\varrho_{A}(\psi)$. The second is expressed in terms of a Markovian approximation to the given state $|\psi\rangle$.
@@ 2 State approximations of decreased complexity
2.1 Reduced state approximation
Given a state vector $|\psi\rangle\in\mathcal{H}_{W}$ and a tripartition $W=A\sqcup B\sqcup C$ in which $B$ is a buffer separating $A$ from the larger region $C$, let $\{|\boldsymbol{\sigma}_{B}\rangle\}$ be an orthonormal basis for $\mathcal{H}_{B}$, e.g. labeled by spin configurations $\boldsymbol{\sigma}_{B}\in\Omega_{B}$ from a convenient computational basis. Let $|\psi(\boldsymbol{\sigma}_{B})\rangle$ denote the normalized vector defined by
$\big{[}\mathbbm{1}_{A}\otimes|\boldsymbol{\sigma}_{B}\rangle\langle\boldsymbol{\sigma}_{B}|\otimes\mathbbm{1}_{C}\big{]}\,|\psi\rangle=:\sqrt{p(\boldsymbol{\sigma}_{B})}\,|\psi(\boldsymbol{\sigma}_{B})\rangle$ (2.1)
with $p(\boldsymbol{\sigma}_{B})\in[0,1]$. In these terms, the state vector is $|\psi\rangle=\sum_{\boldsymbol{\sigma}_{B}}\sqrt{p(\boldsymbol{\sigma}_{B})}\,\,|\psi(\boldsymbol{\sigma}_{B})\rangle$, and its reduced state in $A$ can be presented as
$\varrho_{A}(\psi)=\;\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ \varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)\quad\mbox{with}\quad\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right):=\mathrm{tr}_{\mathcal{H}_{B\sqcup C}}|\psi(\boldsymbol{\sigma}_{B})\rangle\langle\psi(\boldsymbol{\sigma}_{B})|\,.\;$ (2.2)
To interpret this procedure, one may note that the restriction of the reduced state of $|\psi\rangle$ in $A$ is not changed if at first the state is pinched in $B$, that is transformed
$|\psi\rangle\langle\psi|\rightarrow\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\ |\psi(\boldsymbol{\sigma}_{B})\rangle\langle\psi(\boldsymbol{\sigma}_{B})|\,.$ (2.3)
The pinching, which can be viewed as the result of non-destructive measurement of $\boldsymbol{\sigma}_{B}$, allows one to apply to the quantum systems the classical notion of conditioning.
The next observation is that while the rank of $\varrho_{A}(\psi)$ generically reaches the lower of the dimensions $\dim\mathcal{H}_{A}$ and $\dim\mathcal{H}_{BC}$, the density matrices $\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)$ may potentially be approximated by lower rank operators, being conditioned on specified states of the buffer. In particular, an approximation to $\varrho_{A}(\psi)$ of rank bounded by $\dim\mathcal{H}_{B}$ is given by
@@ 2.1 Reduced state approximation
$\boxed{\widehat{\varrho}_{A}\left(\psi\right):=\sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\,|\widehat{\phi}_{A}(\boldsymbol{\sigma}_{B})\rangle\langle\widehat{\phi}_{A}(\boldsymbol{\sigma}_{B})|\,,}$ (2.4)
where ${\widehat{\phi}}_{A}(\boldsymbol{\sigma}_{B})$ is a normalized eigenvector corresponding to the largest eigenvalue, $\lambda(\boldsymbol{\sigma}_{B}):=\left\|\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)\right\|$, of the density matrix in (2.2).
For the fidelity of this approximation, as measured by the trace-norm ($\|\cdot\|_{1}$) distance, one has the bound
$\left\|\varrho_{A}\left(\psi\right)-\widehat{\varrho}_{A}\left(\psi\right)\right\|_{1}\leq 2\ \sum_{\boldsymbol{\sigma}_{B}}p(\boldsymbol{\sigma}_{B})\,\left(1-\lambda(\boldsymbol{\sigma}_{B})\right)\,,$ (2.5)
which follows from straightforward applications of the triangle inequality. Note that $1-\lambda(\boldsymbol{\sigma}_{B})=1-\left\|\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)\right\|$ measures the proximity of $\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)$ to a rank-one operator. If $|\psi(\boldsymbol{\sigma}_{B})\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{BC}$ factorizes into a product state, then $\lambda(\boldsymbol{\sigma}_{B})=1$. One may also explore more general approximations of the reduced state $\varrho_{A}\left(\psi\right)$ through sums over more than one of the leading terms in the spectral decomposition of $\varrho_{A}\left(\psi(\boldsymbol{\sigma}_{B})\right)$.
2.2 Markovian state approximation
For another approach, one may start from a convenient computational orthonormal basis over $\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}$, indexed by spin configurations $(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})\in\Omega_{A}\times\Omega_{B}\times\Omega_{C}$, which represent the eigenvalues of commuting observables. Their joint probability distribution under a given state vector $|\psi\rangle$ is $p(\boldsymbol{\sigma})=|\psi(\boldsymbol{\sigma})|^{2}$ which can be written as
$p(\boldsymbol{\sigma})=p(\boldsymbol{\sigma}_{B})\,p(\boldsymbol{\sigma}_{C}|\boldsymbol{\sigma}_{B})\,p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})$ (2.6)
in terms of the induced conditional probabilities $p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}|\boldsymbol{\sigma}_{B})=p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})/p(\boldsymbol{\sigma}_{B})$ and $p(\boldsymbol{\sigma}_{A}|\boldsymbol{\sigma}_{B})=\sum_{\boldsymbol{\sigma}_{C}}p(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{C}|\boldsymbol{\sigma}_{B})$. Seeking a Markovian approximation of $|\psi\rangle$ it is natural to consider states $\sum_{\boldsymbol{\sigma}}\sqrt{\widehat{p}(\boldsymbol{\sigma})}\,e^{i\widehat{\theta}(\boldsymbol{\sigma})}|\boldsymbol{\sigma}\rangle$ with the modified probability distribution
$\mkern-100.0mu\mathrm{(I)}\mkern 80.0mu\widehat{p}(\boldsymbol{\sigma}_{A},\boldsymbol{\sigma}_{B},\boldsymbol{\sigma}_{C})=p(\boldsymbol{\sigma}_{B})\ p(\boldsymbol{\sigma}_{A}\,|\,\boldsymbol{\sigma}_{B})\ p(\boldsymbol{\sigma}_{C}\,|\,\boldsymbol{\sigma}_{B})\,,$ (2.7)
under which $\boldsymbol{\sigma}_{A}$ and $\boldsymbol{\sigma}_{C}$ retain their marginal distribution but are conditionally independent, and phase functions of the form